OpenStudy (anonymous):

A 1.20 m long ramp is inclined at 20 degree to the horizontal. A solid ball is released from rest at the top of the ramp. Find the ball's speed at the bottom of the ramp. Find its translational acceleration.

OpenStudy (anonymous):

i tried 1.20*sin20=0.410 then 2(9.8)*(.410)=v^2 =2.83 m/s but this is not right what am i doing wrong for the first part? i need help for the first n second part please help

OpenStudy (amistre64):

 gravity is altered by equating g with the the sin sin(t) = g/a a = g csc(t) should be a modified acceleration due to gravity

OpenStudy (amistre64):

now if i use t for time instead of theta ... a(t) = -9.8 csc(20) v(t) = -9.8 csc(20) t s(t) = -4.9 csc(20) + Ho since the ramp is h tall and 1.2 long; that gives us Ho = 1.2 sin(20) as a starting point

OpenStudy (amistre64):

s(t) = -4.9 csc(20) t^2+ 1.2sin(20) this reaches the bottom when t^2 = 1.2sin(20)/4.9csc(20)

OpenStudy (amistre64):

with any luck, thats when t = .1892

OpenStudy (amistre64):

v(.1892) = -9.8csc(20)*(.1892) = abt -5.422 which is speed in a downward direction if i did it right

OpenStudy (amistre64):

i got no idea what a translational accel is tho

OpenStudy (amistre64):

unless that is what i determined to begin with, but without knowing the definition its up for grabs

OpenStudy (anonymous):

so whys g negative?

OpenStudy (anonymous):

* why is

OpenStudy (amistre64):

g is negative so that it pulls the object down the ramp

OpenStudy (amistre64):

its just shows direction of force

OpenStudy (anonymous):

so would that mean that my answer is a negative # as well?

OpenStudy (amistre64):

dunno, im pretty much taking a blind stab at this. if i were to take guess at what they want, i would say that they want the positive value to indicate a magnitude only.

OpenStudy (apoorvk):

i 'd say this question involves rotational dynamics, that is concepts of rolling. (also since its given that the body is a sphere and a solid sphere at that). so, i believe we have to assume that the sphere performs pure rolling and then solve.

OpenStudy (amistre64):

i was wondering in the back of my mind if rolling and "frictionless" where similar